# Leave-Group-Out Cross-Validation for Latent Gaussian Models

### v5

@@ 1.6 Plan of paper
Our paper’s main contribution is the introduction of the automatic LGOCV – a method that intuitively adapts to a specified model. Complementing this, we introduce a computational method to approximate $u_{\text{LGOCV}}$ without needing model refitting. This paper’s computational technique also facilitates the calculation of $u_{\text{LGOCV}}$ with designed groups.
Section 2 introduces LGMs and explains how they can be inferred using INLA. In Section 3, we discuss the automatic group construction method for LGMs. This method can be implemented in two ways: by using the prior correlation matrix or the posterior correlation matrix of the latent linear predictors. In Section 4, we demonstrate how to approximate the LGOCV predictive density. Finally, in Section 5, we compare the approximated LGOCV with the exact LGOCV computed by Markov chain Monte Carlo (MCMC) and present some applications. We conclude with a general discussion in Section 6.
@@ 2 Latent Gaussian models
This section briefly introduces LGMs, as detailed in [8, 9, 10, 11], since the automatic group construction and fast approximation rely on them. The LGMs can be formulated by
$\begin{split}&y_{i}|\eta_{i},\boldsymbol{\theta}\sim\pi(y_{i}|\eta_{i},\boldsymbol{\theta}),\\ &\boldsymbol{\eta}=\boldsymbol{A}\boldsymbol{f},\quad\boldsymbol{f}|\boldsymbol{\theta}\sim N(0,\boldsymbol{P}_{\boldsymbol{f}}(\boldsymbol{\theta})),\boldsymbol{\theta}\sim\pi(\boldsymbol{\theta}).\end{split}$ (2)
@@
3 Automatic Group Construction
The primary feature of automatic group construction is that it requires only a single mandatory parameter to adjust the difficulty of the prediction task. All other processes occur automatically. Our automatic construction intends to make the training set provide less information to predict the testing point such that the LGOCV mimics a task other than interpolation, in which the observed data is assumed to provide less information to predict the new data point. In a multivariate Gaussian distribution, we can quantify the information provided by a data point to predict another data point by the variance reduction of the conditional distribution, and the variance reduction is a function of their correlation coefficient. To elaborate, if $X$ and $Y$ are both Gaussian random variables, the variance reduction resulting from knowing $X$ when predicting $Y$ equates to $\sigma^{2}_{Y}\rho^{2}$, where $\sigma^{2}_{Y}$ is the marginal variance of $Y$ and $\rho$ is the correlation between $X$ and $Y$.
@@ 3 Automatic Group Construction
Manually constructed groups are often based on prior knowledge and some structured effects, represented by $\boldsymbol{f}$. To imitate this process, we can compute the correlation matrix from a submatrix of $\boldsymbol{P}_{\boldsymbol{f}}(\boldsymbol{\theta})$. The correlation matrix, $\boldsymbol{R}_{\text{prior}}$, derived from the submatrix of the prior precision matrix, is a correlation matrix conditioning on those unselected effects. When we are unknowledgeable about our models, we recommend using $\boldsymbol{R}_{\text{post}}$ to construct groups because data will be informative to determine the importance of each effect.
When using a correlation matrix $\boldsymbol{R}$, it is natural to select a fixed number of $\eta_{j}$ most correlated to $\eta_{i}$ and include their index in the group $I_{i}$. However, this approach can be problematic as some linear predictors may have identical absolute correlations to $\eta_{i}$, e.g., in a model with only intercept, all the linear predictors are correlated to each other with correlation $1$. Instead, we include all indices of $\eta_{j}$’s with identical absolute correlations to $\eta_{i}$ in $I_{i}$ if one of them is included. We define a level set as all $\eta_{j}$’s with the same absolute correlation to $\eta_{i}$ and determine the group size based on the number of level sets, denoted as $m$. Setting a higher $m$ results in a less dependent training set and testing point. We recommend using a small number of level sets, such as $m=3$, as a high value of $m$ can result in a large group size.
The automatic group construction process involves selecting the number of level sets, $m$, and the correlation matrix to use, $\boldsymbol{R}_{\text{prior}}$ or $\boldsymbol{R}_{\text{post}}$. For each $i$, we can associate $m$ level sets with the $m$ largest absolute correlations to $\eta_{i}$, and the union of those level sets forms $I_{i}$. As an illustration, we outline the automatic group selection procedure in Algorithm 1.
1 Input: A correlation matrix for latent variables $\boldsymbol{R}$, which is $\boldsymbol{R}_{\text{post}}$ in default, Number of level sets $m$;
2 Output: A list contains groups for all data points;
@@ 4 Approximation of LGOCV predictive density
The integral (6) is computed by the numerical integration [8], and the integral (7) is computed by Gauss-Hermite quadratures [26] as $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ will be approximated by a Gaussian distribution. The key to the accuracy of (7) is that the likelihood, $\pi(y_{i}|\eta_{i},\boldsymbol{\theta})$, is known such that small approximation errors of $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ diminish due to the integration. The accuracy of (6) relies on the accuracy of (7) and the assumption that the removal of $\boldsymbol{y}_{I_{i}}$ does not have a dramatic impact on the posterior.
@@
where $\boldsymbol{\eta}^{*}_{I_{i}}$ is the mode of $\pi_{G}(\boldsymbol{\eta}^{*}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y})$. Note that the correction of hyperparameter reuses $\pi_{G}(\boldsymbol{\eta}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ and $\pi_{G}(\boldsymbol{\eta}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y})$.
@@ 5 Simulations and Applications
This section showcases two simulated examples and two real data applications. We initiate with a simulation that tests the approximation accuracy in a multilevel model with various response types. Following this, a time series forecasting simulation is presented. This allows LGOCV results with automatically constructed groups to be compared to the LFOCV. We then delve into disease mapping, contrasting group constructions derived from various strategies. Finally, we apply our methodology to intricate models using a large dataset, as documented by [27]. The procedures detailed in sections 3 and 4 have been integrated into R-INLA, ensuring that all computational tasks in this section are executed through R-INLA.
@@
This example is a simulation that demonstrates the accuracy of the approximation described in Section 4. The main purpose is to compare $\pi(y_{i}|\boldsymbol{y}_{-I_{i}})$ computed using an approximation in Section 4 and the same quantity computed using MCMC. Additionally, we use the automatic group construction with the number of level sets equal to $1$, corresponding to predicting a data point from a new class.
We simulate data according to the following process. Initially, we simulate 10 class means, denoted as $\boldsymbol{s}$, from a standard normal distribution. Next, we compute $100$ linear predictors, $\eta_{i}=\mu+s_{j(i)}$, where $\mu=\log(10)$ and $j(i)$ is a function mapping data index $i$ to the group index $j$. For this function, we set $j(i)=\lceil\frac{i}{10}\rceil$, where the ceiling function, $\lceil x\rceil$, rounds a number up to the nearest integer. We generate responses according to the linear predictor and response type. The mean of the Gaussian response is $\eta_{i}$, and the standard deviation is 0.1. We generate binomial responses with a probability of $\frac{1}{1+e^{-\eta_{j}}}$ and 20 trials. The exponential responses are generated with a mean of $e^{\eta_{j}}$.
@@
$\begin{split}&\log(\tau_{s})\sim N(0,10^{-4}),\quad\mu\sim N(0,10^{-4}),\quad s_{j}|\tau_{s}\sim N(0,\tau_{s}),\\ &\eta_{i}=\mu+s_{j},\quad y_{i}|\eta_{i}\sim\text{Likelihood}(\eta_{i}),\end{split}$ (12)
where the second parameter of the Gaussian distribution is the precision, and the likelihood is specified according to the data generation process.
@@
11112222333344445555666677778888999910101010001111222233334444555566667777Number of level setsSteps ahead(b) Correspondence between LGOCV and LFOCV
@@
In this real-world example, we will demonstrate the scalability and adaptability of the automatic LGOCV method in a complex model structure and a large sample size. The automatically constructed groups are consistent with the domain knowledge that dengue disease is prevalent in summer.
We will repeat the variable selection process as shown in [27] using the automatic LGOCV. The model chosen by LGOCV is considered to have better predictive power than those selected based on other criteria because the most informative data points for predicting the target are excluded from the training set.
@@
The likelihood of the model is chosen to be negative binomial to account for overdispersion. The latent field consists of a temporal component describing a state-specific seasonality using a cyclic first difference prior distribution and a spatial component describing year-specific spatially unstructured and structured random effects using a modified Besag-York-Mollie (BYM2) model with a scaled spatial component [33]. The temporal component has replications for each state, and the spatial component has replications for each year. We can express the base model using the INLA-style formula,
@@
The results of model selection using deviance information criterion (DIC), LOOCV, and LGOCV are presented in Table 1. The candidate models are those referenced in [32]. To transform equation 1 into a loss function, we calculated its negative value. It is worth noting that LOOCV values across various models exhibit high similarities, suggesting that these models’ interpolation capabilities are similar. This observation is reasonable because all models incorporate spatial and temporal effects that typically offer more predictive power than the covariates. The results obtained from LGOCV are particularly different because of the removal of informative data. This leads to weaker predictive power of the spatial and temporal effects. The importance of the covariates is highlighted as a result.
@@
1 $\text{y}\sim 1+f_{t}+f_{s}$ 3615.38 0.0151 0.0206
2 $\text{y}\sim 1+T_{min}+f_{t}+f_{s}$ 1562.96 0.0064 0.0088
3 $\text{y}\sim 1+T_{max}+f_{t}+f_{s}$ 2228.73 0.0091 0.0133
4 $\text{y}\sim 1+PDSI+f_{t}+f_{s}$ 2167.12 0.0092 0.0126
5 $\text{y}\sim 1+PDSI+T_{min}+f_{t}+f_{s}$ 160.43 0.0006 0.0012
6 $\text{y}\sim 1+PDSI+T_{max}+f_{t}+f_{s}$ 900.65 0.0038 0.0057
7 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{1}+u+f_{t}+f_{s}$ 38.21 0.0002 0*
8 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{2}+u+f_{t}+f_{s}$ 39.13 0.0002 0*
9 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{3}+u+f_{t}+f_{s}$ 28.64 0.0002 0*
10 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{1}+w+f_{t}+f_{s}$ 6.68 0* 0.0005
11 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{2}+w+f_{t}+f_{s}$ 0* 0* 0.0005
12 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{3}+w+f_{t}+f_{s}$ 4.55 0* 0.0006
Note: We offset DIC by 826841.66, LOOCV by 3.2721 and LGOCV by 3.3763.
@@
Figure 5: Groups for testing points from a specific month. $y$-axis: relative frequency, $x$-axis: month of data measurement in groups. The first half-year data are more informative for prediction. As shown in (c) and (d), even in July and November, the group often includes data consistent with the known summer prevalence of dengue.
@@ 6 Discussion
An over-reliance on LOOCV persists in statistical practice, despite concerns raised in studies like [4, 34]. To address this, we have introduced an automatic LGOCV designed to be convenient across various models and to provide a measure of predictive performance. Alongside this, we present an efficient approximation method for computation applicable to manually designed groups. We aim to steer statistical practice towards a better evaluation of predictive performance. It is pertinent to note that our automatic LGOCV do not replace a custom CV strategy designed by modelers tailored for specific applications. Yet, offering an alternative default strategy that improves LOOCV is a significant advantage.

### v6

@@ 1.5 Plan of paper
We propose the model-based LGOCV to evaluate longer-range prediction performance for latent Gaussian models, as a special case of a hierarchical model. Complementing this, we introduce a computational method to approximate $u_{\text{LGOCV}}$ without model refitting, which is crucial for practical implementation of our proposal. Our computational technique also facilitates the calculation of $u_{\text{LGOCV}}$ with user-specified groups.
Section 2 introduces LGMs and explains how they can be efficiently inferred using INLA. In Section 3, we discuss the model-based group construction method for LGMs. This method can be implemented in two ways: by using the prior correlation matrix or the posterior correlation matrix of the latent linear predictors. In Section 4, we demonstrate how to approximate the LGOCV predictive density. Finally, in Section 5, we compare the approximated LGOCV with the exact LGOCV computed by Markov chain Monte Carlo (MCMC) and present some applications. We conclude with a general discussion in Section 6.
@@ 2 Latent Gaussian models
This section briefly introduces LGMs, as detailed in [20, 21, 22, 23], since the model-based group construction and fast approximations rely on them. The LGMs can be formulated by
$\begin{split}&y_{i}|\eta_{i},\boldsymbol{\theta}\sim\pi(y_{i}|\eta_{i},\boldsymbol{\theta}),\\ &\boldsymbol{\eta}=\boldsymbol{A}\boldsymbol{f},\quad\boldsymbol{f}|\boldsymbol{\theta}\sim N(0,\boldsymbol{P}_{\boldsymbol{f}}(\boldsymbol{\theta})),\quad\boldsymbol{\theta}\sim\pi(\boldsymbol{\theta}).\end{split}$ (2)
@@
3 Model-based Group Construction
The primary feature of our proposed group construction is that it requires a choice of correlation matrix (prior or posterior) for the linear predictors, and a single mandatory parameter to adjust the difficulty of the prediction task. This parameter is termed "the number of level sets". It can be interpreted as the strength of the non-dependence between the group to leave out and the rest of the data. A higher value would thus ensure that the leave-out group is more independent from the rest of the data, than a lower value. A higher indpeendnece between the leave-out data and the rest of the data simluates a more difficult prediction task for the model. Based on this value and the correlation matrix choice, all other processes are automated. In a multivariate Gaussian distribution, we can quantify the information provided by a data point to predict another data point by the variance reduction of the conditional distribution, and the variance reduction is a function of their correlation coefficient. To elaborate, if $X$ and $Y$ are both Gaussian random variables, the variance reduction resulting from knowing $X$ when predicting $Y$ equates to $\sigma^{2}_{Y}\rho^{2}$, where $\sigma^{2}_{Y}$ is the marginal variance of $Y$ and $\rho$ is the correlation between $X$ and $Y$.
@@ 3 Model-based Group Construction
Manually constructed groups are often based on prior knowledge and some structured effects, represented by $\boldsymbol{f}$. To imitate this process, we can compute the correlation matrix from a submatrix of $\boldsymbol{P}_{\boldsymbol{f}}(\boldsymbol{\theta})$. The correlation matrix, $\boldsymbol{R}_{\text{prior}}$, derived from the submatrix of the prior precision matrix, is a correlation matrix conditioning on those unselected effects. The groups constructed using $\boldsymbol{R}_{\text{prior}}$ are viewed to be solely user-defined in the way that it only depends on the priors and not on the data. In some situations this could be motivated, but in general we recommend using $\boldsymbol{R}_{\text{post}}$ to construct groups because the data will be informative to determine the importance of each effect.
When using a correlation matrix $\boldsymbol{R}$, it is natural to select a fixed number of $\eta_{j}$ most correlated to $\eta_{i}$ and include their index in the group $I_{i}$. However, this approach can be problematic as some linear predictors may have identical absolute correlations to $\eta_{i}$, e.g., in a model with only intercept, all the linear predictors are correlated to each other with correlation $1$. Instead, we include all indices of $\eta_{j}$’s with identical absolute correlations to $\eta_{i}$ in $I_{i}$ if one of them is included. We define a level set as all $\eta_{j}$’s with the same absolute correlation to $\eta_{i}$ and determine the group size based on the number of level sets, denoted as $m$. Setting a higher $m$ results in a less dependent training set and testing point. We recommend using a small number of level sets, such as $m=3$, as a high value of $m$ can result in a large leave-out group size.
The automated group construction process thus involves selecting the number of level sets, $m$, and the correlation matrix to use, $\boldsymbol{R}_{\text{prior}}$ or $\boldsymbol{R}_{\text{post}}$. For each $i$, we can associate $m$ level sets with the $m$ largest absolute correlations to $\eta_{i}$, and the union of those level sets forms $I_{i}$. As an illustration, we outline the automated group construction procedure in Algorithm 1.
1 Input: A correlation matrix choice $\boldsymbol{R}$ (posterior correlation is the default), Number of level sets $m$;
2 Output: A list containing the groups for all data points;
3 Calculate $\boldsymbol{R}$ from the model;
@@ 4 Approximation of LGOCV predictive density
The integral (6) is computed by the numerical integration [20], and the integral (7) is computed by Gauss-Hermite quadratures [33] as the conditional posterior density $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ will be approximated by a Gaussian distribution. The key to the accuracy of (7) is that the likelihood, $\pi(y_{i}|\eta_{i},\boldsymbol{\theta})$, is known such that small approximation errors of $\pi(\eta_{i}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ diminish due to the integration. The accuracy of (6) relies on the accuracy of (7) and the assumption that the removal of $\boldsymbol{y}_{I_{i}}$ does not have a dramatic impact on the posterior.
@@
where $\boldsymbol{\eta}^{*}_{I_{i}}$ is the mode of $\pi_{G}(\boldsymbol{\eta}^{*}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y})$. Note that the correction of the hyperparameter reuses $\pi_{G}(\boldsymbol{\eta}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y}_{-I_{i}})$ and $\pi_{G}(\boldsymbol{\eta}_{I_{i}}|\boldsymbol{\theta},\boldsymbol{y})$.
@@ 5 Simulations and Applications
This section showcases two simulated examples and two real data applications. The code is available at https://github.com/zhedongliu/LGOCV.
We start with a simulation that tests the approximation accuracy in a multilevel model with various response types. Following this, a time series forecasting simulation is presented. This allows LGOCV results with automatically constructed groups to be compared to the LFOCV. We then delve into disease mapping, contrasting group constructions derived from various strategies. Finally, we apply our methodology to intricate models using a large dataset, as documented by [34]. For the construction of the leave-out group for LGOCV, we used Algorithm 1 in Section 3. The procedures detailed in Sections 3 and 4 have been integrated into R-INLA, ensuring that all computational tasks in this section are executed through R-INLA.
@@
This example is a simulation that demonstrates the accuracy of the approximation described in Section 4. The main purpose is to compare $\pi(y_{i}|\boldsymbol{y}_{-I_{i}})$ computed using an approximation in Section 4 and the same quantity computed using MCMC. Furthermore, we use automatic group construction with the number of level sets equal to $1$, corresponding to predicting a data point from a new class.
We simulate data according to the following process. Initially, we simulate 10 class means, denoted as $\boldsymbol{s}$, from a standard normal distribution. Next, we compute $100$ linear predictors, $\eta_{i}=\mu+s_{j(i)}$, where $\mu=\log(10)$ and $j(i)$ is a function mapping data index $i$ to the group index $j$. For this function, we set $j(i)=\lceil\frac{i}{10}\rceil$, where the ceiling function, $\lceil x\rceil$, rounds a number up to the nearest integer. We generate responses according to the linear predictor and one of three response types; Gaussian, Binomial and Exponential. The mean of the Gaussian response is $\eta_{i}$, and the standard deviation is 0.1. We generate binomial responses with a success probability of $\frac{1}{1+e^{-\eta_{j}}}$ for 20 trials. The exponential responses are generated with a mean of $e^{\eta_{j}}$.
@@
$\begin{split}&\log(\tau_{s})\sim N(0,10^{-4}),\quad\mu\sim N(0,10^{-4}),\quad s_{j}|\tau_{s}\sim N(0,\tau_{s}),\\ &\eta_{i}=\mu+s_{j},\quad y_{i}|\eta_{i}\sim\text{response model}(\eta_{i}),\end{split}$ (12)
where the second parameter of the Gaussian distribution is the precision, and the likelihood is specified according to the data generation process with the given response model.
@@
111122223333444455556666777788889999101010100011112222333344445555666677778888999910101010Number of level setsSteps ahead(b) Correspondence between LGOCV and LFOCV
@@
In this real-world example, we will demonstrate the scalability and adaptability of the automatic LGOCV method in a complex model structure and a large sample size. The automatically constructed model-based groups are consistent with the domain knowledge that dengue disease is prevalent in summer.
We will repeat the variable selection process as shown in [34] using the automatic LGOCV. The model chosen by LGOCV is considered to have better predictive power for longer-range predictions, than those selected based on other criteria because the most informative data points for predicting the target are excluded from the training set.
@@
The data generating model is chosen to be negative binomial, to account for overdispersion. The latent field consists of a temporal component describing a state-specific seasonality using a cyclic first difference prior distribution and a spatial component describing year-specific spatially unstructured and structured random effects using a modified Besag-York-Mollie (BYM2) model with a scaled spatial component [40]. The temporal component has replications for each state, and the spatial component has replications for each year. We can express the base model using the INLA-style formula,
@@
The results of model selection using deviance information criterion (DIC), LOOCV, and LGOCV ($m=2,3,4$) are presented in Table 1. The candidate models are those referenced in [39]. To transform equation 1 into a loss function, we calculated its negative value.
From Table 1 we note that LOOCV prefers the spatio-temporal model that incorporates access to water while the spatio-temporal model with urbanization as a covariate, is preferred by LGOCV. This result is interesting since we can conclude that the same model might not necessarily perform well for short- and longer-range prediction. The practitioner thus needs to decide what the goal of the modeling is, and then choose the model to be used accordingly. If we want to predict dengue risk for a new unobserved area or time point, it seems that urbanization has a better prediction ability than access to water. Note also that as we increase the number of level sets, we are defining a prediction target with an increased range, thus moving further away from LOOCV.
@@
($m=2$) ($m=3$) ($m=4$)
1 $\text{y}\sim 1+f_{t}+f_{s}$ 3615.38 0.0151 0.0158 0.0206 0.0270
2 $\text{y}\sim 1+T_{min}+f_{t}+f_{s}$ 1562.96 0.0064 0.0067 0.0088 0.0098
3 $\text{y}\sim 1+T_{max}+f_{t}+f_{s}$ 2228.73 0.0091 0.0098 0.0133 0.0163
4 $\text{y}\sim 1+PDSI+f_{t}+f_{s}$ 2167.12 0.0092 0.0095 0.0126 0.0184
5 $\text{y}\sim 1+PDSI+T_{min}+f_{t}+f_{s}$ 160.43 0.0006 0.0006 0.0012 0.0023
6 $\text{y}\sim 1+PDSI+T_{max}+f_{t}+f_{s}$ 900.65 0.0038 0.0038 0.0057 0.0084
7 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{1}+u+f_{t}+f_{s}$ 38.21 0.0002 0* 0* 0*
8 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{2}+u+f_{t}+f_{s}$ 39.13 0.0002 0* 0* 0*
9 $\text{y}\sim 1+PDSI+T_{min}+PDSI*u_{3}+u+f_{t}+f_{s}$ 28.64 0.0002 0* 0* 0*
10 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{1}+w+f_{t}+f_{s}$ 6.68 0* 0.0005 0* 0.0014
11 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{2}+w+f_{t}+f_{s}$ 0* 0* 0.0005 0* 0.0015
12 $\text{y}\sim 1+PDSI+T_{min}+PDSI*w_{3}+w+f_{t}+f_{s}$ 4.55 0* 0.0006 0* 0.0014
@@
Note: We offset DIC by 826841.66, LOOCV by 3.2721, LGOCV ($m=2$) by 3.314, LGOCV ($m=3$) by 3.3763 and LGOCV ($m=4$) by 3.4372.
@@
Figure 5: Groups for testing points from a specific month. $y$-axis: relative frequency, $x$-axis: month of data measurement in groups. The first half-year data are more informative for prediction. As shown in (c) and (d), even in July and November, the group often includes data consistent with the known summer prevalence of dengue. Note that dengue is prevalent in the summer months which are approximately November to February.
@@ 6 Discussion
An over-reliance on LOOCV to evaluate predictive capacity in general persists in statistical practice, despite concerns raised in studies such as [17, 41]. LOOCV can provide an evaluation of short-range predictive ability with well-established asymptotics for some models. On the other hand, what can we do to evaluate the longer-range predictive ability of complex models? Various approaches for specific models, such as time series or spatial models have been proposed, where custom CV procedures are designed to mimic a longer-range prediction task than that of LOOCV. We have introduced an automated approach for evaluating the longer-range prediction ability of any latent Gaussian model, namely LGOCV. LGOCV is designed to be applicable to all models that are latent Gaussian models, and thus provides a framework for general longer-range predictive ability evaluation without the need for case-by-case considerations. Moreover, we propose a computationally efficient approach to calculate LGOCV scores and metrics based on the INLA methodology. We have shown that our approximate LGOCV implementation is almost exact when compared with the results from MCMC, albeit at a much lower computational cost. This enables practitioners to use the LGOCV approach for complex models and large data.
Our approach is designed for latent Gaussian models and some ideas can thus be extended to the case of non-latent Gausian models with careful consideration of the computational cost associated with this endeavor. LGOCV for LGMs is computationally efficient in INLA, since it is fully parallelizable by computing the necessary quantities only at the mode of the hyperparameters. For huge data ($n>10^{6}$) however, the cost will be high since the cost increases linearly in $n$, albeit much lower than other available approaches. For huge data, performing CV on a subset or constructing the groups manually could be considered. Nonetheless, for LGMs, the proposed LGOCV could be considered as the most feasible approach for longer-range predictive ability evaluation.
The choice of the number of level sets determine the prediction task and thus the degree of independence between the leave-out group and the rest of the data. There is not a one-to-one correspondence between the number of level sets and the number of points to leave out as shown in the simulations and applications, although a higher number of level sets would imply a longer range for the prediction task, than a lower number. The choice of the number of level sets remains arbitrary since it is a user-defined parameter, we recommend a low number like $m=2$ or $m=3$ if there is no clear indication of what else $m$ should be. There exists no optimal value of $m$ in general, since it would imply different prediction tasks for different levels of dependency. In our applications, and those of others who have applied the LGOCV framework, it is shown that LGOCV provides the information we need to evaluate longer-range prediction ability, and complements the information from LOOCV.
It is pertinent to note that the proposed LGOCV do not replace a custom CV strategy designed by modelers, tailored for specific applications. We pose it as an alternative default strategy for longer-range prediction ability evaluation, that complements LOOCV in assessing the predictive ability of an LGM, while being computationally efficient and practical for real-world scenarios.